A solution of hydrochloric acid in water is 60.0% HCl and 40% water, H2O, by weight. The solution has a density of 1.19 g/mL. What is the


b. Molarity of the water (with hydrochloric acid as the solvent)

c. Molality of water in hydrochloric acid?

d. mole fraction of hydrochloric acid?

e. mole fraction of water?


1
Expert's answer
2022-03-23T09:13:03-0400

b. Molarity of water (with hydrochloric acid as the solvent)

100.0 g of the solution contains 60.0 g of HCl and 40.0 g of H2O

Moles HCl: 60.0 / 36.5 = 1.64 mol

Moles H2O: 40.0 / 18.0 = 2.22 mol

Volume of the solution: d = m/V; V = m/d = 100.0 / 1.19 = 84.0 (mL or 0.084 L)

Molarity of water: 2.22 / 0.084 = 26.4 (M)  

c. Molality of water in hydrochloric acid?

    Molality of water = 2.22 mol / 0.06 kg = 37 mol/kg

d. mole fraction of hydrochloric acid?

    Mole fraction HCl = 1.64 / (1.64 + 2.22) = 0.42 or 42%

e. mole fraction of water?

    Mole fraction water = 100% - 42% = 58%.